OpenStudy (mikeymaximum):

Replace ? with =, >, or < to make the statement true. 16 ÷ 4 + 12 ? 4 + 24 ÷ 2 A. > B. < C. =

OpenStudy (mikeymaximum):

I got C.

OpenStudy (anonymous):

Left Side =16/4+12 =4+12 =16 Right Side 4+24/2 =4+12 = 16 16=16 so C.